This one confuses me. Its from the A+ program.
Sean earns $4.50 per hour more than his trainee, Tyrone. During one 8 hour day, they earn a total of $284. How much does Sean make per hour?
What I know so far is:
S = Sean T = Tyrone
T + $4.50 = S
(S + T)8 = $284

Answer by checkley71(8403)

let x=houly rate of the trainee.
8x+8(x+4.50)=284
8x+8x+36=284
16x=284-36
16x=248
x=248/16
x=$15.50 is the hourly salary for the trainee.
15.50+4.50=$20 for Sean.
proof
8*15.50+8*20=284
124+160=284
284=284